A baseball player hits a ball straight up into the air. If the 0.142-kg baseball reached a maximum height of 32 m above the ground, how much gravitational potential energy did it gain by being hit? Assume that the baseball was 0.98 m above the ground when it was hit.
1
Expert's answer
2018-04-26T10:22:08-0400
The law of conservation of energy states
E_i=E_f
E_hit+mgh_i=mgh_f
So
E_hit=mg(h_f-h_i)
=0.142×9.81×(32.00-0.98)=43.2 J